# Sharp Time Decay and Regularization

### v1

@@ Abstract
We consider a class of porous medium type of equations with Caputo time derivative. The prototype problem reads as $D^{\alpha}_{t}u=-\mathcal{L}u^{m}$ and is posed on a bounded Euclidean domain $\Omega\subset\mathbb{R}^{N}$ with smooth boundary and zero Dirichlet boundary conditions. The linear operator $\mathcal{L}$ falls within a wide class of either local or nonlocal operators, and the nonlinearity is allowed to be of degenerate or singular type, namely, $10<m<$ 1 and $m>1$.
Using the fractional gradient flow theory, we show existence of unique solution and prove new $L^{p}-L^{\infty}$ smoothing effects by employing the Green function method. The comparison principle, which we provide in the most general setting, serves as a crucial tool in the proof and provides a novel monotonicity formula. To the best of our knowledge, this type of smoothing effects are novel in nonlinear PDEs involving Caputo derivative. Consequently, we establish that the regularizing effect due to the diffusion is stronger than the memory effect introduced by the fractional time derivative. We also prove that the solution satisfies the zero Dirichlet condition pointwise by means of sharp boundary estimates.
Finally, we prove that our solution does not vanish in finite time if $10<m<$ 1, unlike the case with the classical time derivative. Indeed, we provide a sharp rate of decay for any $L^{p}$-norm of the solution for any $m>0$. Our findings indicate that the difference between slow and fast diffusion is mitigated by the memory effect, which dramatically slows the spatial diffusion.
Keywords: porous medium equation; fast diffusion equation; nonlocal operators; Caputo fractional time derivative; subdiffusion; comparison principle; regularity estimates; long time behavior. MSC2020 Classification. Primary: 34K37, 35K55, 35B51, 35B65, 35R11, 35B40. Secondary: 35K61, 35K65, 35K67.
@@ Acknowledgments
M.B. and P.I. where partially supported by the Project PID2020-113596GB-I00 (Ministry of Science and Innovation, Spain) and by the E.U. H2020 MSCA program, grant agreement 777822. M.B. acknowledges financial support from the Spanish Ministry of Science and Innovation, through the “Severo Ochoa Programme for Centres of Excellence in R&D” (CEX2019-000904-S). P.I. was partially funded by the FPU-grant FPU19/04791 from the Spanish Ministry of Universities. This work started when P.I. was visiting M.G. at The University of Texas at Austin in the year 2022, when an essential part of the work was done. P.I. would like to thank both M.G. and the Mathematical Department of The University of Texas at Austin for the kind hospitality, and the above MSCA project for the financial support.
M.G. is partially supported by NSF grants DMS-2206677 and DMS-1514761. M.G. also thanks the Isaac Newton Institute (Cambridge, UK) and ICMAT (Madrid, Spain) for their kind hospitality.
@@ 1 Introduction
We consider the Dirichlet problem for a general class of Porous Medium type of equations with Caputo fractional time derivative of order $\alpha\in(0,1)$:
@@
with $u(x,t)$ defined on a smooth bounded domain $\Omega\subset\mathbb{R}^{N}$. The linear operator $\mathcal{L}$ belongs to a general class of densely defined operators both of local and nonlocal type, which includes the classical Laplacian and fractional Laplacian operator. The fractional time operator $D_{t}^{\alpha}$ is the Caputo derivative, introduced in [23] and defined for $\alpha\in(0,1)$ by
$D^{\alpha}_{t}f(t)=\frac{1}{\Gamma(1-\alpha)}\frac{\rm d}{\,{\rm d}t}\int_{0}^{t}\frac{f(\tau)-f(0)}{(t-\tau)^{\alpha}}\,{\rm d}\tau\,.$ (1)
In the last years, subdiffusive partial differential equations – equations that combine diffusion operators in the space variable with fractional time derivatives – have attracted significant attention, as they are used to model anomalous diffusion, viscoelastic phenomena and processes with memory [40, 41, 42]. The prototype subdiffusive PDE is the “Heat Equation with memory”,
$D^{\alpha}_{t}u=\Delta u\,.$ (2)
For the well-posedness and regularity of (2) in $\mathbb{R}^{N}$, see for instance the works of Allen, Caffarelli and Vasseur in [3]; Kemppainen, Siljander, Vergara and Zacher in [34]; and Zacher in [52]. Optimal asymptotic decay for solutions of (2) are obtained by Cortazar, Quiros and Wolansky in [24] for the Cauchy problem in $\mathbb{R}^{N}$. For the Dirichlet problem on bounded domains, Vergara and Zacher [49] show long time decay estimates, also allowing for general operators with variable coefficients in space and time. Recently, Chan, Gomez-Castro and Vázquez studied in [25] global well-posedness of (CPME) with $m=1$ for singular solutions on bounded domains.
@@
Beside global well-posedness, the main contributions of this manuscript are threefold: (i) the study of comparison principle and time monotonicity formula, (ii) $L^{p}-L^{\infty}$ smoothing effects, and (iii) optimal long time behaviour for (CPME). A surprising fact established with our analysis is the non-extinction in finite time for all solutions of (CPME) when $m\in(0,1)$. This fact was previously observed in [49] for separate-variable solutions. In this work, we prove this property for any general solution.
The Caputo derivative adds the extra difficulty of the “memory effect”. Roughly speaking, this memory effect somehow destroys the semigroup structure, essential in the De Giorgi-Nash-Moser theory. Surprisingly enough, a nontrivial adaptation of the so-called Green function method, introduced in [17], allows us to establish our main results, overcoming the memory effect. This is achieved by employing novel time monotonicity estimates, a feature that comes as a surprise in the context of the Caputo setting, as we shall further comment below.
To provide further insights about our results, some numerology is in order: our asymptotic estimates correspond to the known ones in the formal limit $m\rightarrow 1$. On the other hand, this does not happen in the –formal– limit $\alpha\rightarrow 1^{-}$. Indeed, neither the exponents of smoothing effects nor the ones in the long time behaviour estimates, correspond to the known ones when $\alpha=1$.
@@
Comparison principle and time monotonicity. We first recall the results about existence and uniqueness of solution for (CPME) from [1, 37]. One of the approximations steps proposed in [37] is the starting point of our proof of the comparison principle. The comparison principle here corresponds to “well-ordering”: if initial data for (CPME) are ordered, then solutions will preserve the same order for any positive time. In the present case, we have to overcome two difficulties: we do not have the standard semigroup property $S(t+h)f=S(h)S(t)f$ because of the Caputo derivative, and we lack regularity to perform rigorously the computations. For these reasons, our proof involves several approximation steps. This comparison principle of Theorem 1.2 lays the cornerstone for several other results in this manuscript, most importantly for a new time monotonicity estimates “a la Benilan-Crandall”. Namely, we show that $t^{\frac{\alpha}{m-1}}u(t)$ is monotone in time for every $t>0$, cf. Theorem 1.3. The unexpected nature of this result arises from the memory effect, which complicates any form of time monotonicity.
Smoothing effects and boundary estimates. We show new $L^{p}-L^{\infty}$ smoothing effects, stated in Theorem 1.4. To the best of our knowledge, these kind of $L^{p}-L^{\infty}$ regularizing effects have been proved by Kemppainen, Siljander, Vergara and Zacher in [34] only for the linear Caputo Heat equation (2) on the whole space, whenever the initial data is integrable enough. Our regularizing estimates hold for $u_{0}\in L^{p}(\Omega)$ with $p>\frac{N}{2s}$. There is a structural reason why one should not expect similar results for $p<\frac{N}{2s}$: the fundamental solution of (2) in $\mathbb{R}^{N}$ is unbounded and belongs to $L^{q}(\mathbb{R}^{N})$ with $q<\frac{N}{N-2}$. Therefore, one should ask the initial datum to be in $L^{p}(\mathbb{R}^{N})$ with $p>\frac{N}{2}$.
@@
We also study the behaviour of solutions near the boundary and show sharp estimates of the form $u(t,x)\lesssim{\rm dist}^{\gamma/m}(x,\partial\Omega)t^{-\alpha/m}$ in Theorem 1.5. Indeed, for “big initial data” we obtain lower estimates with matching powers, proving the so-called Global Harnack Principle, that is $u(t,x)\asymp{\rm dist}^{\gamma/m}(x,\partial\Omega)t^{-\alpha/m}$, see Theorem 1.6. These results are optimal both in space and in time, the latter optimality is explained in what follows.
@@
Optimal long time behaviour. Our last contribution concerns a sharp analysis of the long time behaviour of the solution, cf. Theorem 1.7: in contrast to the local-in-time results, in the fast diffusion regime $m\in(0,1)$, solutions to (CPME) do not extinguish in finite time. This is an important side effect of the “memory” intrinsic in the Caputo derivative. We compute the rates of decay of all $L^{p}$ norms from above and below, with matching powers. This phenomenon of non-extinction was first observed by Vergara and Zacher in [49] for separate-variables solutions. Here, by means of differential inequalities involving Caputo derivative [^1] We devote Appendix A to recall the main results used here about ODE with Caputo derivative., we show that the particular behaviour of separate-variables solutions is enjoyed by all the other solutions as well, and is optimal. We cover the whole range $m>0$ and a wide class of operators $\mathcal{L}$, see Section 1.2 and 1.3 for more details about the diffusion operators. This aspect reveals a remarkable discontinuity in the parameter $\alpha$: indeed we show here that solutions decay as $u(t)\asymp t^{-\alpha/m}$, while in the local-in-time case $\alpha=1$ they are known to decay as $u(t)\asymp t^{-1/(m-1)}$ when $m>1$ and extinguish in finite time $u(t)\asymp(T-t)_{+}^{1/(1-m)}$ when $m\in(0,1)$.
@@
Related results. We conclude this section by recalling previous work on the porous medium (PME) and fast diffusion equations (FDE) with the classical time derivative. The literature on these models is extremely rich and spans several decades. The few references we mention here are by no means exhaustive. Both the porous medium and fast diffusion equations have been widely used to describe different physical phenomena, including flow of gas through porous media [46], plasma in reactors [8], etc. For this reason and for their own mathematical interest, they have attracted a lot of attention since the early 50s, see for instance the lecture notes of Aronson [5] and the books of Barenblatt [7], Vázquez [45, 46] and Daskalopulos and Kenig [26]. See also the papers [44] ($m>1$) and [9] ($m<1$).
@@
In the nonlocal case, we refer to the survey [47] for a complete review up to 2016, and we just mention the works strictly related to ours. When dealing with the Cauchy problem for the nonlocal PME and FDE, namely, $u_{t}=-(-\Delta)^{s}u^{m}$ on the whole space, a complete theory has been developed by De Pablo, Quirós, Rodríguez and Vázquez in [28, 29, 30, 31]; see also [16, 48]. As for the Dirichlet problem in the degenerate case (i.e. $\alpha=1$ and $m>1$ in the present paper), Figalli, Ros-Oton, Sire, Vazquez and the first author, performed a sharp analysis of the behaviour of the solutions in a series of papers [17, 15, 10, 12, 11, 18]: existence and uniqueness (in optimal classes), comparison, smoothing effects, infinite speed of propagation (contrary to the local case $\mathcal{L}=-\Delta$), local and global Harnack inequalities, sharp interior and boundary regularity estimates, sharp asymptotic behaviour. When $m<1$ the theory is still under development: existence, uniqueness, smoothing effects and boundary estimates have been recently proven by Ispizua, the first and last authors in [13], while Harnack and higher regularity estimates shall appear in [14]. The proofs rely on the so-called Green function method, introduced in [17], which offers a valid alternative to the more-classical approaches “a la” De Giorgi-Nash-Moser; see [13, Section 3] for a comparison between the two methods.

### v3

# Time-Fractional Porous Medium Type Equations:

@@ Abstract
We consider a class of porous medium type of equations with Caputo time derivative. The prototype problem reads as $D^{\alpha}_{t}u=-\mathcal{L}u^{m}$ and is posed on a bounded Euclidean domain $\Omega\subset\mathbb{R}^{N}$ with zero Dirichlet boundary conditions. The operator $\mathcal{L}$ falls within a wide class of either local or nonlocal operators, and the nonlinearity is allowed to be of degenerate or singular type, namely, $10<m<$ 1 and $m>1$. This equation is the most general form of a variety of models used to describe anomalous diffusion processes with memory effects, and finds application in various fields, including visco-elastic materials, signal processing, biological systems and geophysical science.
We show existence of unique solution and new $L^{p}-L^{\infty}$ smoothing effects. The comparison principle, which we provide in the most general setting, serves as a crucial tool in the proof and provides a novel monotonicity formula. Consequently, we establish that the regularizing effects from the diffusion are stronger than the memory effects introduced by the fractional time derivative. Moreover, the solution attains the boundary conditions pointwise.
Finally, we prove that the solution does not vanish in finite time if $10<m<$ 1, unlike the case with the classical time derivative. Indeed, we provide a sharp rate of decay for any $L^{p}$-norm of the solution for any $m>0$. Our findings indicate that memory effects weaken the spatial diffusion and mitigate the difference between slow and fast diffusion.
Keywords: porous medium equation; fast diffusion equation; nonlocal operators; Caputo fractional time derivative; subdiffusion; comparison principle; regularity estimates; long time behavior.
MSC2020 Classification. Primary: 34K37, 35K55, 35B51, 35B65, 35R11, 35B40. Secondary: 35K61, 35K65, 35K67.
@@ Acknowledgments
M.B. and P.I. were partially supported by the Project PID2020-113596GB-I00 (Ministry of Science and Innovation, Spain) and by the E.U. H2020 MSCA program, grant agreement 777822. M.B. acknowledges financial support from the Spanish Ministry of Science and Innovation, through the “Severo Ochoa Programme for Centres of Excellence in R&D” (CEX2019-000904-S). P.I. was partially funded by the FPU-grant FPU19/04791 from the Spanish Ministry of Universities.
This work started in Fall 2022 when P.I. was visiting M.G. at The University of Texas at Austin, where an essential part of the work was done. P.I. would like to thank both M.G. and the Mathematical Department of The University of Texas at Austin for the kind hospitality, and the above MSCA project for the financial support.
M.G. is partially supported by NSF grants DMS-2206677 and DMS-1514761. M.G. also thanks the institute ICMAT (Madrid, Spain) for the kind hospitality.
@@ 1 Introduction
Several phenomena, from physics to biology to finance, exhibit events during which fractional behavior and memory effects become predominant; for example viscoelastic materials (whose response depends on their current and past states), certain geographical processes including movement of groundwater or transportation through porous media, neuronal and gene regulation networks, but also control theory and more recent modeling of the financial market. Fractional calculus provides a reliable tools to describe memory effects. In 1967 M. Caputo, in the context of modeling heterogeneous elastic fields, introduced a novel differential operator,
$D^{\alpha}_{t}f(t)=\frac{1}{\Gamma(1-\alpha)}\frac{\rm d}{\,{\rm d}t}\int_{0}^{t}\frac{f(\tau)-f(0)}{(t-\tau)^{\alpha}}\,{\rm d}\tau,\,\quad\alpha\in(0,1).$ (1)
Operator (1) took later the name of Caputo time derivative of order $\alpha$. In subsequent research, Caputo modeled certain type of fluid diffusing in porous media using this novel nonlocal operator, [28, 29, 30]. In these geothermal studies, Darcy’s Law is adapted to describe fluids that may carry solid particles obstructing the pores, thus diminishing their size and creating a pattern of mineralization. This phenomenon has recently been observed in various other types of porous materials, including building materials [44, 45, 64] and zeolite [37, 38].
Systems where particles exhibit anomalous diffusion (sub- or super-diffusive behavior) often involve memory effects [58, 59, 66]. Examples include diffusion in porous media, turbulent flows, and biological transport processes. These applications require mathematical models that allow particles to do macroscopical long jumps (Lévy flights), leading to the use of nonlocal operators in the spatial and time variables [5, 48, 58] . Many different mathematical models describing anomalous diffusion in a porous medium can be found in the literature. The one we consider here reads as
@@
and includes a general class of densely defined operators $\mathcal{L}$ both of local and nonlocal type. Equation (CPME) is a density dependent diffusion resulting in a characteristic scaling $x/t^{\alpha/(2+N(m-1))}$, whenever the diffusion operator is $\mathcal{L}=-\Delta$. Its derivation can be found in [63] and some of its applications can be read here [37, 38, 44, 45, 64]. We briefly recall its derivation in Section 1.2.
From a mathematical point of view, the characteristic scaling of the wetting front variable in anomalous diffusion differs from the one of the classical Heat Equation ($x/\sqrt{t}$). Originally, Caputo derivative arose in the linear setting to achieve a subdiffusive characteristic scaling of the form $x/t^{\alpha/2}$ with $\alpha\in(0,1)$. The non-locality in time, or the memory effect, represents a “waiting time” phenomenon typically derived within the stochastic framework of the Continuous Time Random Walk, see for instance [58, 60].
In view of the above discussion, (CPME) encompasses a wide variety of anomalous diffusion models that combine local and nonlocal spatial operators, Caputo fractional time derivative, and $m$-power like nonlinearities for any $m>0$. In this manuscript, we provide a comprehensive qualitative and quantitative study of this model.
Beside global well-posedness, the main contributions of this manuscript are threefold: (i) study of comparison principle and time monotonicity formula, (ii) $L^{p}-L^{\infty}$ smoothing effects, and (iii) optimal long time behaviour. The surprising facts established with our analysis are the regularizing effects and the non-extinction in finite time for all solutions of (CPME) when $m\in(0,1)$. Notably, the memory effect slows down the diffusion, minimizing the relevance of the nonlinearity parameter $m$ in the ranges $m\in(0,1)$ and $m>1$. Nevertheless, the diffusive nature of the equation still provides a regularization of the solution, to the best of our knowledge, a feature previously unknown for nonlinear equations involving Caputo derivatives. We emphasize that all these results are new even for the classical Laplacian $\mathcal{L}=-\Delta$ and $\alpha\in(0,1)$.
Within the mathematical framework, the prototype subdiffusive PDE is the “Heat Equation with memory”,
$D^{\alpha}_{t}u=\Delta u\,,$ (2)
and the literature on this topic is vaste. For the well-posedness and regularity of (2) in $\mathbb{R}^{N}$, see for instance the works of Allen, Caffarelli and Vasseur in [3]; Kemppainen, Siljander, Vergara and Zacher in [50]; and Zacher in [77]. Optimal asymptotic decay for solutions of (2) are obtained by Cortazar, Quiros and Wolansky in [31] for the Cauchy problem in $\mathbb{R}^{N}$. For the Dirichlet problem on bounded domains, Vergara and Zacher [73] show long time decay estimates, also allowing for general operators with variable coefficients in space and time. Recently, Chan, Gomez-Castro and Vázquez studied in [33] global well-posedness of (CPME) with $m=1$ for singular solutions on bounded domains.
@@
The memory effects complicate the analysis quite a lot. Roughly speaking, the memory effects somehow destroy the semigroup structure, essential in the De Giorgi-Nash-Moser theory. Surprisingly enough, a nontrivial adaptation of the so-called Green function method, introduced in [20], allows us to establish our main results. This is achieved by employing novel time monotonicity estimates, a feature that comes as a surprise in the context of the Caputo setting, as we shall further comment below.
To provide further insights about our results, some numerology is in order: our asymptotic estimates correspond to the known estimates in the formal limit $m\rightarrow 1$. On the other hand, this does not happen in the –formal– limit $\alpha\rightarrow 1^{-}$. Indeed, neither the exponents of smoothing effects nor the ones in the long time behaviour estimates, correspond to the known ones when $\alpha=1$.
@@
Comparison principle and time monotonicity. We first recall the results about existence and uniqueness of solution for (CPME) from [1, 54]. One of the approximations steps proposed in [54] is the starting point of our proof of the comparison principle. The comparison principle here corresponds to “well-ordering”: if initial data for (CPME) are ordered, then solutions will preserve the same order for any positive time.
In the present case, we have to overcome two difficulties: we do not have the standard semigroup property $S(t+h)f=S(h)S(t)f$ because of the Caputo derivative, and we lack regularity to perform rigorously the computations. For these reasons, our proof involves several approximation steps. It is remarkable that we inherit comparison properties from the discretized problem, in such a general framework (see more comments on that after Theorem 1.2). This comparison principle lays the cornerstone for several other results, most importantly for a new time monotonicity estimates “a la Benilan-Crandall”. Namely, we show that $t^{\frac{\alpha}{m-1}}u(t)$ is monotone in time for every $t>0$, cf. Theorem 1.3. The unexpected nature of this result arises from the memory effect, which complicates any form of time monotonicity.
Smoothing effects and boundary estimates. We show new $L^{p}-L^{\infty}$ smoothing effects, stated in Theorem 1.4. To the best of our knowledge, these kind of $L^{p}-L^{\infty}$ regularizing effects have been proved by Kemppainen, Siljander, Vergara and Zacher in [50] only for the linear Caputo Heat equation (2) on the whole space, whenever the initial data is integrable enough. Our regularizing estimates hold for $u_{0}\in L^{p}(\Omega)$ with $p>\frac{N}{2}$. There is a structural reason why one should not expect similar results for $p<\frac{N}{2}$: the fundamental solution of (2) in $\mathbb{R}^{N}$ is unbounded and belongs to $L^{q}(\mathbb{R}^{N})$ with $q<\frac{N}{N-2}$. Therefore, one should ask the initial datum to be in $L^{p}(\mathbb{R}^{N})$ with $p>\frac{N}{2}$ being the conjugate exponent of $q$.
@@
Lastly, we study the behaviour of solutions near the boundary and show sharp estimates of the form $u(t,x)\lesssim{\rm dist}^{\gamma/m}(x,\partial\Omega)t^{-\alpha/m}$ in Theorem 1.5 for some $\gamma\in(0,1]$. Indeed, for “big initial data” we obtain lower estimates with matching powers, proving the so-called Global Harnack Principle, that is $u(t,x)\asymp{\rm dist}^{\gamma/m}(x,\partial\Omega)t^{-\alpha/m}$, see Theorem 1.6. These results are optimal both in space and in time, the latter optimality is explained in what follows.
The case of unbounded domains. All the above results can be carried out, essentially with similar proofs, to the case of (CPME) posed on unbounded domains, including the whole space $\mathbb{R}^{N}$. See Remark 1 for more details.
@@
Optimal long time behaviour. Our last contribution concerns a sharp analysis of the long time behaviour, cf. Theorem 1.7: in contrast to the local-in-time results, in the fast diffusion regime $m\in(0,1)$, solutions to (CPME) do not extinguish in finite time. This is an important side effect of the “memory” intrinsic in the Caputo derivative. We compute the rates of decay of all $L^{p}$ norms from above and below, with matching powers. This phenomenon of non-extinction was first observed by Vergara and Zacher in [73] for separate-variables solutions. Here, by means of differential inequalities involving Caputo derivative [^1] We devote Appendix A to recall the main results used here about ODE with Caputo derivative., we show that the particular behaviour of separate-variables solutions is enjoyed by all the other solutions as well, and is optimal. This aspect reveals a remarkable discontinuity in the parameter $\alpha$: indeed we show here that solutions decay as $u(t)\asymp t^{-\alpha/m}$, while in the local-in-time case $\alpha=1$ they are known to decay as $u(t)\asymp t^{-1/(m-1)}$ when $m>1$ and extinguish in finite time $u(t)\asymp(T-t)_{+}^{1/(1-m)}$ when $m\in(0,1)$.
@@
Related results. We conclude this section by recalling previous work on the porous medium (PME) and fast diffusion equations (FDE) with the classical time derivative. The literature on these models is extremely rich and spans several decades. The few references we mention here are by no means exhaustive. Both the porous medium and fast diffusion equations have been widely used to describe different physical phenomena, including flow of gas through porous media [70], plasma in reactors [9], etc. For this reason and for their own mathematical interest, they have attracted a lot of attention since the early 50s, see for instance the lecture notes of Aronson [6] and the books of Barenblatt [8], Vázquez [69, 70] and Daskalopulos and Kenig [35]. See also the papers [7, 68] ($m>1$) and [11] ($m<1$).
@@
In the nonlocal case, we refer to the survey [71] for a complete review up to 2016, and we just mention the works strictly related to ours. When dealing with the Cauchy problem for $u_{t}=-(-\Delta)^{s}u^{m}$, a complete theory has been developed by De Pablo, Quirós, Rodríguez and Vázquez in [39, 40, 41, 42]; see also [19, 72]. As for the Dirichlet problem in the degenerate case (i.e. $\alpha=1$ and $m>1$ in the present paper), Figalli, Ros-Oton, Sire, Vazquez and the first author, performed a sharp analysis of the behaviour of the solutions in a series of papers [12, 13, 14, 18, 20, 21]: existence and uniqueness (in optimal classes), comparison, smoothing effects, infinite speed of propagation (contrary to the local case $\mathcal{L}=-\Delta$), local and global Harnack inequalities, sharp interior and boundary regularity estimates, sharp asymptotic behaviour. When $m<1$ the theory is still under development: existence, uniqueness, smoothing effects and boundary estimates have been recently proven by Ispizua, the first and last authors in [16], while Harnack and higher regularity estimates shall appear in [17]. The proofs rely on the so-called Green function method, introduced in [20], which offers a valid alternative to the more-classical approaches “a la” De Giorgi-Nash-Moser; see [16, Section 3] for a comparison between the two methods.